How does the increase of wavelength by 50% affect the frequency of a wave that travels in a rope?


1
Expert's answer
2021-02-19T18:57:49-0500

The relationship between the frequency and the wavelength of the wave can be written as follows:


"f\lambda=v."

From this formula we can find the frequency of the wave:


"f=\dfrac{v}{\lambda}."

Let the wavelength increases by 50%, then we can write:


"f_{new}=\dfrac{v}{1.5\lambda}=\dfrac{f}{1.5}=0.67f."

As we can see from the calculations, the new frequency is 67% of the old frequency. In other words, the old frequency decreases by 33%.

Answer:

If the wavelength of the wave increases by 50% the frequency of the wave decreases by 33%.